A child's playground swing is supported by chains that are 4.00 m long. If a child in the swing is 0.63 m above the ground and moving at 5.20 m/s when the chains are vertical, what is the maximum height of the swing ? Assume the masses of the chains are negligible.
1
Expert's answer
2021-05-25T16:40:18-0400

"y_i =" initial height of swing "= 0.63m"


"v_i =" initial speed of swing "= 5.20m\/s"


We have to find the maximum height "y_{max}" of the swing


Applying the principle of conservation of energy


"\Delta K+ \Delta U = 0"


"(\dfrac{1}{2}mv_f^2-\dfrac{1}{2}mv_i^2)+mg\Delta y = 0"


where "v_f = 0" at the maximum height of the swing:


"-\dfrac{1}{2}mv_i^2+(mgy_{max}-mgy_i) = 0"


"mgy_{max} = mgy_i+\dfrac{1}{2}mv_i^2"


Hence,

"y_{max} = \dfrac{gy_i+\dfrac{1}{2}v_i^2}{g}"

"y_{max} = \dfrac{(9.8)(0.63)+\dfrac{1}{2}(5.20)}{9.8}"


"y_{max} = 0.89m"